insert a rational number and irrational number between:
(i)2 and 3(ii)1/3and 1/2

Answer by stanbon(75887)

insert a rational number and irrational number between:
Comment: The average of the 2 numbers is inbetween.
------------------
(i)2 and 3
x = (2+3)/2 = 2.5
----
(ii)1/3 and 1/2
y = [(1/3)+(1/2)]/2
y = (5/6)/2
y = 5/12
====================
cheers,
Stan H.